# GRPO-LEAD: A Difficulty-Aware Reinforcement Learning Approach for Concise Mathematical Reasoning in Language Models

### v1

@@ 4.2 Hyperparameters
We train with a learning rate of $1\times 10^{-6}$, batch size 32, and group size 8–generating 8 rollouts per question for GRPO reward computation. The KL penalty term is removed, as it was found to suppress exploration in our experiments, which is also suggested in similar worksLiu et al. (2025); Hu et al. (2025).
For the length-dependent accuracy reward, we set $\alpha=0.05$, providing a moderate decay that encourages conciseness without penalizing slight verbosity.
For difficulty-aware advantage reweighting, we use $A=0.4$, $B=1.5$, $\rho_{0}=0.75$, and $k=10$. This configuration ensures reweighting is minimal on easy problems but sharply increases near the 75% correctness threshold. The steep slope ($k=10$) enables strong emphasis on high-difficulty examples, guiding the model to allocate learning more effectively.
@@ 4.3 Model Variants and Fine-Tuning Stages
Starting from the DeepSeek-R1 Distilled 7B Qwen-Math checkpoint, we first apply standard GRPO on the 9k questions, producing a baseline. Then, we train 3 more models from the DeepSeek-R1 Distilled 7B Qwen-Math checkpoint, adding one more of the following components subsequently: (i) Length Reward only, (ii) Length Reward + Advantage Reweighting, (iii) Length Reward + Advantage Reweighting + Explicit Penalty. We train for approximately 200 steps and select the top-performing checkpoints based on validation results. At test time, we limit the generation length to 8k for all 7B models, matching the training length limit.
@@
We extend the above procedure to the DeepSeek-R1 Distilled 14B Qwen checkpoint across multiple stages. In Stage 1, we train for 100 steps using all GRPO-LEAD components on the filtered 9k-question dataset. To evaluate the benefit of supervised fine-tuning (SFT), we first fine-tune the model on a curated set of 13k math problems, then apply GRPO. This SFT stage significantly improves the model’s reasoning quality, even though it tends to increase the output length and caused some format error.
The supervised fine-tuning data consists of all problems in the DeepScaler dataset with difficulty greater than 1. To construct high-quality reasoning traces for SFT, we use the QWQ-32B modelTeam (2025) to generate step-by-step solutions.
After observing that some questions remain low correctness, we further fine-tune for Stage 2 with an additional 100 steps, focusing on these underperformed problems. Lastly, in Stage 3, we address repetitive output patterns by removing the length penalty and introducing a negative reward ($-1.5$) for repeated $n$-grams. We continue training for 140 more steps, yielding the final model checkpoint. At test time, we limit the generation length to 14k for all 14B models, in accordance with our training settings and also to better evaluate the models’ performance in a low-budget scenario.
@@ 4.4 Baselines and Evaluation Protocol
Table 1: Ablation results on AIME24 and AIME25. We report Cons@32 (the fraction of problems for which at least one correct solution is found among 32 samples), Pass@1, and the average token length (Lenavg). The best value in each column is in boldface, the second best is underlined.
@@
Deepseek-7B 0.767 0.431 6990 0.467 0.292 7113
GRPO + len. reward 0.767 0.438 5275 0.533 0.308 5210
     + adv. reweighting 0.767 0.458 5323 0.567 0.325 5437
     + explicit penalty 0.800 0.470 6104 0.567 0.345 6308
Model Name AIME24 AIME25
Cons@32 Pass@1 Lenavg Cons@32 Pass@1 Lenavg
DeepSeek-14B 0.800 0.614 9182 0.633 0.429 10046
Light-R1-14B-DS 0.833 0.641 9571 0.767 0.505 10194
LEAD-stage1 0.833 0.629 8790 0.767 0.523 9371
LEAD-stage3 0.867 0.650 8267 0.767 0.539 8668
5 Results
@@ 5.2 Ablation Analysis
Effect of Length Reward
We first incorporate the length-dependent accuracy reward into GRPO. Compared to Deepseek-7B, on AIME24, this maintains the original Cons@32 (0.767) while slightly improving Pass@1 from 0.431 to 0.438. Notably, the average solution length substantially decreases from 6990 to 5275 tokens, a reduction of approximately 24.5%. Similarly, on AIME25, Cons@32 improves from 0.467 to 0.533, accompanied by a significant length reduction of nearly 1900 tokens (approximately 26.8%). These results demonstrate, length reward, by penalizing correct but overly verbose solutions, can effectively reduce unnecessary text without compromising overall performance.
Effect of Advantage Reweighting
Adding difficulty-aware advantage reweighting further refines performance. On AIME24, although Cons@32 remains 0.767, Pass@1 increases to 0.458. On AIME25, both Cons@32 and Pass@1 improve (0.533 $\rightarrow$ 0.567 and 0.308 $\rightarrow$ 0.325, respectively). These results demonstrate that prioritizing challenging problems strengthens the model’s reasoning robustness, as the reweighting strategy mitigates over-reliance on simpler examples. This validates our hypothesis that calibrating training focus toward harder instances drives more reliable generalization.
Effect of Explicit Penalty for Incorrect Answers
Finally, we introduce a negative reward term to penalize incorrect solutions explicitly. As shown in Table 1, this addition yields the highest Pass@1 scores across the board (0.470 on AIME24 and 0.345 on AIME25). Cons@32 also climbs to 0.800 on AIME24. Note, however, that solution length increases modestly from about 5300 tokens to 6104 on AIME24, reflecting a trade-off: while the explicit penalty effectively sharpens the decision boundary and boosts accuracy, the model also tends to be more conservative and invests more tokens to ensure correctness. Nonetheless, the resulting average solution length is still lower than the baseline Deepseek-7B.
@@ 5.3 Comparison with Baselines
We next evaluate GRPO-LEAD at the 14B scale and compare it against two strong baselines under a 14k-token generation budget: DeepSeek-14B and the state-of-the-art Light-R1-14B-DS. Table 2 presents results on AIME24 and AIME25, including both our intermediate model (LEAD-stage1) and our final model (LEAD-stage3).
@@
LEAD-stage1 achieves a Cons@32 of 0.833, matching Light-R1-14B-DS and exceeding DeepSeek-14B by 4.1% (0.833 vs. 0.800). Its Pass@1 (0.629) outperforms DeepSeek-14B (0.614) and closely approaches Light-R1-14B-DS (0.641). Crucially, LEAD-stage1 produces more concise responses (8790 tokens) than both baselines. Building on these gains, LEAD-stage3 pushes performance further, delivering the highest Cons@32 (0.867, 4% above Light-R1-14B-DS) and the best Pass@1 (0.650), while reducing average solution length to 8267 tokens.
@@
LEAD-stage1 yields a Cons@32 of 0.767, on par with Light-R1-14B-DS (0.767) and substantially ahead of DeepSeek-14B (0.633). Its Pass@1 (0.523) outperforms both DeepSeek-14B (0.429) and Light-R1-14B-DS (0.505). Once again, solutions from LEAD-stage1 are notably shorter (9371 tokens) than those of the baselines. In turn, LEAD-stage3 attains the highest Cons@32 (0.767) and Pass@1 (0.539), while further trimming average length to 8668 tokens.
Overall, both LEAD-stage1 and LEAD-stage3 deliver substantial improvements over DeepSeek-14B and Light-R1-14B-DS, simultaneously boosting correctness and conciseness under a constrained (14k-token) budget. Remarkably, training LEAD-stage1 for just 100 steps—requiring only about 24 hours on eight H20 GPUs—already matches Light-R1-14B-DS on Cons@32 and outperforms it on AIME25 Pass@1 while producing shorter solutions, underscoring the practical efficiency of GRPO-LEAD for large-scale math problem-solving.
6 Limitations
Additionally, we only have access to a limited amount of compute, which prevents us from conducting more comprehensive experiments. For instance, we currently cannot provide the validation curve for the 7B model in the ablation study that adds an explicit penalty. This is due to an error in the validation code after upgrading to the newest VERL version, and we currently don’t have the compute to reproduce it. We also couldn’t formally perform a hyperparameter search to showcase the rationale behind choosing the hyperparameters for our designed modifications.

### v2

@@ 4.2 Hyperparameters
We train with a learning rate of $1\times 10^{-6}$, batch size 32, and group size 8—generating 8 rollouts per question for GRPO reward computation. The KL penalty term is removed, as it was found to suppress exploration in our experiments, which is also suggested in similar works Liu et al. (2025b); Hu et al. (2025b).
For the length-dependent accuracy reward, we set $\alpha=0.05$, providing a moderate decay that encourages conciseness without penalizing slight verbosity. For difficulty-aware advantage reweighting, we use $A=0.4$, $B=1.5$, $\rho_{0}=0.75$, and $k=10$. This configuration ensures reweighting is minimal on easy problems but increases sharply near the 75% correctness threshold. The steep slope ($k=10$) enables strong emphasis on high-difficulty examples, guiding the model to allocate learning more effectively.
@@ 4.3 Model Variants and Fine-Tuning Stages
Starting from the DeepSeek-R1 Distilled 7B Qwen-Math checkpoint, we first apply standard GRPO on the 9k questions, producing a baseline. Then, we train three more models from the DeepSeek-R1 Distilled 7B Qwen-Math checkpoint, adding one more of the following components subsequently: (i) Length Reward only, (ii) Length Reward + Advantage Reweighting, (iii) Length Reward + Advantage Reweighting + Explicit Penalty. We train for approximately 200 steps and select the top-performing checkpoints based on validation results. At test time, we limit the generation length to 8k for all 7B models, matching the training length limit.
@@
We extend the above procedure to the DeepSeek-R1 Distilled 14B Qwen checkpoint across multiple stages. In Stage 1, we train for 100 steps using all GRPO-LEAD components on the filtered 9k-question dataset. To enhance the model’s base capability, we first fine-tune the model on a curated set of 13k math problems with supervised fine-tuning (SFT), then conduct the RL phase. This SFT stage significantly improves the model’s reasoning quality, even though it tends to increase the output length and caused some format errors.
The SFT data consists of all problems in the DeepScaler dataset with difficulty greater than 1. To construct high-quality reasoning traces for SFT, we use the QWQ-32B model Team (2025) to generate step-by-step solutions.
After observing that some questions remain low correctness, we further fine-tune for Stage 2 to focus on those underperformed problems. We also address the repetitive output patterns by removing the length penalty and introducing a negative reward (1.5-1.5) for repeated $n$-grams. We continue training for 240 more steps (100 steps with initial settings and 140 more steps with repetition penalty), yielding the final model checkpoint. At test time, we limit the generation length to 14k for all 14B models, in accordance with our training settings and also to better compare the models’ performance in a low-budget scenario.
@@ 4.4 Baselines and Evaluation Protocol
5 Results
@@
Deepseek-7B 0.767 0.431 6,990 0.467 0.292 7,113
GRPO + len. reward 0.767 0.438 5,275 0.533 0.308 5,210
     + adv. reweighting 0.767 0.458 5,323 0.567 0.325 5,437
     + explicit penalty 0.800 0.470 6,104 0.567 0.345 6,308
Table 1: Ablation results on AIME24 and AIME25. We report Cons@32 (accuracy through majority voting for 32 samplings), Pass@1, and the average token length (Lenavg). The best value in each column is in boldface, the second best is underlined.
@@ 5.1 Training Dynamics
Figure 2: Validation Pass@1 over training steps for three configurations: GRPO, GRPO+L, and GRPO+LAD. As shown by the faster convergence, length reward and advantage reweighting provide a richer reward signal signal than the original setup.
11footnotetext: The validation consists of 27 challenging problems from AIMO2 Frieder et al. (2024), CMU-MATH-AIMO Sun (2024), and AIME24.
@@ 5.2 Ablation Analysis
(a) AIME24
(b) AIME25
Figure 3: Performance against inference budget for training done with different ablations of LEAD. GRPO with length reward (GRPO+L) largely enhances the performance at low budget settings compared to before training (DeepseekR1-7B).
@@
Length Reward Brings Conciseness to Reasoning
We first incorporate the length-dependent accuracy reward into GRPO. Compared to Deepseek-7B, length reward slightly improves Pass@1 on both AIME24 by 1.6% ($0.431\rightarrow 0.438$) and AIME25 by 5.4% ($0.292\rightarrow 0.308$), with an additional improvement of Cons@32 by 14.1% on AIME25. Notably, these improvements are accompanied by a substantial reduction of 1,715 tokens (24.5%) and 1,903 tokens (26.8%) in the average response length on the two datasets, respectively. Figure 3 further demonstrates that length reward largely enhances performance in low-budget settings over the base model, matching its peak performance with only 5/8 of the token budget on the more difficult AIME25. These results demonstrate that length reward, by penalizing correct but overly verbose solutions, can effectively reduce unnecessary text without compromising overall performance.
Advantage Reweighting Encourages Model to Solve More Difficult Problems
Further incorporating difficulty-aware advantage reweighting (GRPO+LAD) refines performance. On AIME24, Pass@1 increases from the GRPO+L stage by 4.8% ($0.438\rightarrow 0.458$), while Cons@32 remains $0.767$. For AIME25, both Pass@1 and Cons@32 improve by 5.5% ($0.308\rightarrow 0.325$) and 6.4% ($0.533\rightarrow 0.567$), respectively. As Figure 3 shows, GRPO+LAD demonstrates gains over GRPO+L in almost all budget regimes on AIME25 and for budgets exceeding 5k tokens on AIME24. These results indicate that advantage reweighting, by prioritizing challenging problems, strengthens reasoning robustness and mitigates over-reliance on simpler examples, thus validating its role in driving more reliable generalization.
Model Name AIME24 AIME25
Cons@32 Pass@1 Lenavg Cons@32 Pass@1 Lenavg
DeepSeek-14B 0.800 0.614 9,182 0.633 0.429 10,046
Light-R1-14B-DS 0.833 0.641 9,571 0.767 0.505 10,194
LEAD-stage1 0.833 0.629 8,790 0.767 0.523 9,371
LEAD-stage2 0.867 0.650 8,267 0.767 0.539 8,668
Explicit Penalty for Incorrect Answers Regularizes Thinking
Finally, introducing an explicit penalty for incorrect solutions (GRPO+LEAD) yields the highest Pass@1 scores. On AIME24, Pass@1 and Cons@32 improve from the GRPO+LAD stage by 2.6% ($0.458\rightarrow 0.470$) and 4.3% ($0.767\rightarrow 0.800$), respectively. On AIME25, Pass@1 also increases by 6.2% ($0.325\rightarrow 0.345$), as detailed in Table 1. Notably, these gains involve a modest increase in average solution length on AIME24 (from approximately 5,300 to 6,104 tokens). Figure 3 illustrates this trade-off, showing a performance sacrifice in low-budget regimes, though GRPO+LEAD still outperforms GRPO+LAD with budgets higher than 5k tokens on AIME25. These results suggest that the explicit penalty serves as a regularizer for the model to be more conservative about its reasoning. Such regularization boosts performance while requiring a slightly longer thinking process, which nevertheless remains shorter than the Deepseek-7B baseline.
@@ 5.3 Comparison with Baselines
We next evaluate GRPO-LEAD at the 14B scale and compare it against two strong baselines under a 14k-token generation budget: DeepSeek-14B and the state-of-the-art Light-R1-14B-DS. Table 2 presents results on AIME24 and AIME25, including both our intermediate model (LEAD-stage1) and our final model (LEAD-stage2).
@@
LEAD-stage1 achieves a Cons@32 of 0.833, matching Light-R1-14B-DS and exceeding DeepSeek-14B by 4.1%. Its Pass@1 outperforms DeepSeek-14B by 2.4% and closely approaches Light-R1-14B-DS. Crucially, LEAD-stage1 produces more concise responses than both baselines, with more than 800 tokens less on average. Building on these gains, LEAD-stage2 pushes performance further, delivering the highest Cons@32 (4% above Light-R1-14B-DS) and the best Pass@1, while reducing average solution length to 8,267 tokens.
@@
LEAD-stage1 yields a Cons@32 of 0.767, matching Light-R1-14B-DS and exceeding DeepSeek-14B by 21.2%. Its Pass@1 (0.523) outperforms DeepSeek-14B by 21.9% and Light-R1-14B-DS by 3.6%. Crucially, LEAD-stage1 produces more concise responses than both baselines, with its solutions averaging 9,371 tokens. Building on these gains, LEAD-stage2 pushes performance further, delivering the highest Cons@32 (matching Light-R1-14B-DS at 0.767) and the best Pass@1 (0.539), while reducing average solution length to 8,668 tokens.
Overall, both LEAD-stage1 and LEAD-stage2 deliver substantial improvements over DeepSeek-14B and Light-R1-14B-DS, simultaneously boosting correctness and conciseness under a constrained (14k-token) budget. Remarkably, training LEAD-stage1 for just 100 steps—requiring only about 24 hours on eight H20 GPUs—already matches Light-R1-14B-DS on Cons@32 and outperforms it on AIME25 Pass@1 while producing shorter solutions, underscoring the practical efficiency of GRPO-LEAD for large-scale math problem-solving.
@@ 6 Conclusion
7 Limitations
Additionally, we only have access to a limited amount of compute, which prevents us from conducting more comprehensive experiments. For instance, we currently cannot provide the validation curve for the 7B model in the ablation study that adds an explicit penalty. This is due to an error in the validation code after upgrading to the newest VERL version, and we currently do not have the compute to reproduce it. A comparison with the original GRPO model is also missing, except for the curve shown in Figure 2, because the checkpoint was stored on a rented server that was automatically released as we were writing the paper. We also couldn’t formally perform a hyperparameter search to showcase the rationale behind choosing the hyperparameters for our designed modifications.
Appendix A Evaluations on Coding Tasks
We evaluate our proposed LEAD-14B model against the original DeepSeek-R1-Distill-Qwen-14B baseline on the LiveCodeBench benchmark under a maximum sequence length of 8k tokens. The dataset version used is release_v5, consisting of 880 code generation tasks. Results are summarized in Table 3.
Model Cons@32 Avg. Correct Avg. Answer Precision Pass@1
Normal Problems (1–5)
Deepseek-7B 0.8 18.8 20.3 0.708 0.588
GRPO + L 0.8 19.7 27.6 0.631 0.616
GRPO + LAD 0.9 20.1 26.9 0.687 0.628
GRPO + LEAD 0.8 22.0 24.5 0.723 0.688
Difficult Problems (6–10)
Deepseek-7B 0.4 8.3 13.8 0.404 0.259
GRPO + L 0.5 8.6 24.1 0.412 0.269
GRPO + LAD 0.6 9.8 24.2 0.448 0.306
GRPO + LEAD 0.6 9.7 20.0 0.421 0.303
Highly Difficult Problems (11–15)
Deepseek-7B 0.2 0.9 2.0 0.230 0.028
GRPO + L 0.3 1.3 13.9 0.172 0.041
GRPO + LAD 0.2 1.3 14.6 0.156 0.041
GRPO + LEAD 0.4 1.5 7.7 0.355 0.047
Table 4: Consolidated evaluation results on AIME25, stratified by problem difficulty. Avg. Answers refers to the number of outputs that have completed within the 8k token budget and produce some, where as Avg. Correct refers to the correct answers. Precision is Avg. Correct/ Avg. Answers.
As shown above, LEAD-14B achieves higher accuracy (0.5156 vs. 0.5103) while producing slightly longer completions. This suggests that our method enhances reasoning capability in code generation. Regarding the observed increase in chain-of-thought (CoT) length, we hypothesize that this effect arises because our training focused exclusively on mathematical reasoning datasets. While our method compresses reasoning paths in math domains, such compression does not appear to generalize as effectively to code. Combined with the improved reasoning capability that may increase the overall reasoning path, this may explain why generated sequences are overall longer in code-related tasks.
Appendix B Detailed Analysis on AIME25 by Difficulty
To further analyze model performance, we stratified the AIME25 dataset into three difficulty levels based on the problem number: normal (problems 1–5), difficult (problems 6–10), and highly difficult (problems 11–15). The detailed evaluation results for each stratum are presented in Table 4.
The stratified results in Table 4 support the hypothesis that advantage reweighting enhances a model’s ability to solve more difficult problems. This is evidenced by the widening performance gap in Pass@1 between GRPO+L and GRPO+LAD as problem difficulty increases. For normal problems, GRPO+LAD offers a modest $1.95\%$ improvement over GRPO+L. This margin increases substantially to $13.7\%$ for difficult problems, indicating that the benefits of advantage reweighting are more pronounced in challenging scenarios.
For highly difficult problems, the Pass@1 scores for GRPO+L and GRPO+LAD are identical. Neither method incorporates an explicit penalty for incorrect answers, making them susceptible to generating numerous wrong solutions. This tendency leads to unstable majority voting-based accuracy (Cons@32), a vulnerability that is magnified by the intrinsic difficulty of the problems.
In contrast, the introduction of an explicit penalty in GRPO+LEAD demonstrates a clear regularization effect. On the most difficult problem set, GRPO+LEAD achieves the highest accuracy (Cons@32 of $0.4$) and more than doubles the precision of both GRPO+L ($0.172$) and GRPO+LAD ($0.156$); the number of correct answers generated by GRPO+LEAD is comparable to both GRPO+L and GRPO+LAD, despite generating much fewer total answers. This validates our hypothesis that the explicit penalty effectively "regularizes thinking", discouraging the kind of hasty and incorrect responses that the length reward tends to encourage otherwise.
Aspect GRPO+L Deepseek-7B
Structure & flow Step 1: enumerate all possible triples … Step 2: compute the multinomial coefficient … Step 3: sum and mod.” “Okay … let me parse this step by stepbut wait, hold onlet me verify the triples again …”
Redundancy “…hence, all possible triples: (6, 2, 1), (5, 3, 1), (4, 3, 2).” “So, the possible triples … So, three triples in totalWait, hold on, let me check if there are moreSo, total three triples.
Conciseness of language “Total N = 2016. Therefore, the remainder is 16.” Wait, hold on a second.Maybe I can think of all possible partitionsNo, I think the only possible triples are the three we found.
Logical signposting Case 1: s = 1 … Case 2: s = 2 … Case 3: s = 3 (no solutions).” Case 1: S = 1 … Subcase 1aSubcase 1b(digression)Case 3: S = 3 … no solutions(returns to earlier cases).
Error-checking “Only three possible triples, so the computation is complete.” Wait, hold on a second. Is that all?let me verify the triples againmaybe there are other triples?
Length Entire solution $\approx$ 200 words. Entire solution $\approx$ 370 words (many repeated sentences such as “So, I think 16 is the answer.”).
Table 5: Qualitative comparison of the shortest correct rollouts from GRPO+L and Deepseek-7B for AIME 25 I, Problem 3. Italicized text in the Deepseek-7B column represents meta-commentary or self-correction loops.
Appendix C Qualitative Analysis of Solution Conciseness
To provide a qualitative illustration of how the length reward enhances conciseness, we contrast the shortest correct solutions generated by GRPO+L and the baseline Deepseek-7B for the same problem (Problem 3, AIME 25 I). Table 5 breaks down the comparison across key aspects of readability and reasoning structure.
As the comparison highlights, the GRPO+L model produces a tight, step-by-step solution that remains focused, avoids repetition, and concludes efficiently. In contrast, the Deepseek-7B baseline’s reasoning path is less direct, characterized by repeated self-checks and conversational digressions that nearly double the total length and reduce clarity. This case study demonstrates that our length-reward mechanism successfully encourages a more disciplined and economical reasoning style.